# Beyond Homophily: Community Search on Heterophilic Graphs

### v1

@@ IV-A Encoder Frameworks
Figure 2 gives an end-to-end overview of our AdaptCS encoder. Starting from an input graph $G=(V,E,X)$, the model framework disentangles spectral frequency and hop distance, then fuses the resulting 2D signal by a channel mixer.
Stage 1: Distinctive-hop channel (Sec.IV-B). Given a node $v_{i}$, we recursively conduct $k$-hop neighborhoods aggregation to construct distance-aware features. The $k$-th channel only aggregates information from neighbors with an exact shortest distance of $k$, which eliminates redundancy and noisy messages during convolution. We then apply the weight-renormalization operator to amplify signals to maintain numerical stability.
Stage 2: Frequency-aware channel (Sec.IV-C). Frequency filters process the hop-specific feature matrices to produce two complementary views: a smooth (homophilic) representation and a non-smooth (heterophilic) representation. This separation allows each channel to specialize in either cohesive similarity or distinctive contrast without mutual interference.
Stage 3: 2D channel mixing (Sec.IV-D). In this stage, we first use attention to fuse the low- and high-frequency channels within each hop, and then concatenate the fused representations across all hops. Next, a multi-layer perceptron (MLP) or a per-class attention bank is applied, producing a unified node embedding that integrates both local details and long-range context in a compact representation.
@@ IV-B Distinctive Hop Channels (Distance-aware)
In this section, we propose distinctive hop channels to explicitly process signals from different hops of neighbors, tailored for heterophilic graphs. Distinctive hop channels address three limitations in traditional multi-hop aggregation approaches. First, this approach solves the flip effect, where signals from each hop are processed independently, handling the complex semantics between high-order relations. Second, the distinctive hop adjacency matrices are shortest-distance aware, meaning each node in the $k$-th hop channel accurately represents edges with the shortest path distance of exactly $k$. Third, this approach reduces redundancy and improves space efficiency since the sum of all distinctive hop adjacency matrices precisely reconstructs the cumulative adjacency $\hat{A}^{k}$, avoiding repetitive edge representation.
Motivation. To overcome the flip effect, which arises when multi-hop aggregation erroneously reverses node similarities, we generalize the diversification distinguishability framework proposed in ACM [ACM_2022] to incorporate higher-order structural information. Specifically, we define the High-Order Diversification Distinguishability (HDD) as follows:
Definition IV.1 (High-Order Diversification Distinguishability (HDD)).
Given a graph $G$ with $c>2$ classes, let $H^{(k)}_{\mathrm{HP}}$ denote the high-pass (heterophily-enhancing) embedding at hop $k$ for each node, and let the fused high-pass embedding be
${H}_{\mathrm{HP}}=\mathcal{F}\big(H^{(1)}_{\mathrm{HP}},\ldots,H^{(k)}_{\mathrm{HP}}\big)$
where $\mathcal{F}(\cdot)$ is a hop fusion operator (e.g., attention, sum, or MLP). Define the similarity between nodes $v$ and $u$ as:
$S_{v,u}=\langle{H}_{\mathrm{HP},v},\ {H}_{\mathrm{HP},u}\rangle,$
where $\langle\cdot,\cdot\rangle$ denotes the inner product or cosine similarity. A node $v$ is high-order diversification distinguishable if:
  1. 1.
    $\mathrm{Mean}_{u}\left(\{S_{v,u}\mid u\in V,\ Z_{u}=Z_{v}\}\right)\geq 0$;
  2. 2.
    $\mathrm{Mean}_{u}\left(\{S_{v,u}\mid u\in V,\ Z_{u}\neq Z_{v}\}\right)\leq 0$.
The HDD value of the graph is defined as: $\mathrm{HDD}_{{H}_{\mathrm{HP}}}(G)=\frac{1}{|V|}\left|\{v\mid v\in V\land v\text{ satisfies HDD conditions}\}\right|.$
The HDD measures whether the aggregated embeddings clearly distinguish same-class neighbors from different-class neighbors at every hop, thus providing a direct indicator of the flip effect problem. Specifically, the flip effect arises because conventional methods merge all hop-distance information indiscriminately, and, in multi-class settings ($c>2$), two nodes of different classes sharing a common heterophilic neighbor can exhibit spurious positive similarity after high-pass filtering. By employing distinctive-hop masking, each channel is restricted to aggregating information distinctively from each hop, which separates negative two-hop effects from direct interactions. The 2D mixer also ensures the fused embedding is free from noise. Consequently, each hop’s embedding captures accurate semantics, eliminating ambiguity and flip effects.
Therefore, under ideal conditions (e.g., one-hot labels), our distinctive-hop filtering ensures the HDD criterion is fully satisfied for every node, leading to the theoretical guarantee:
Theorem 1 (Distance awareness Yields HDD = 1 for $c>2$).
Suppose $X=Z$ (one-hot labels) and $\hat{A}^{(k)}$ is constructed via distinctive-hop masking. For any number of classes $c>2$, all nodes become high-order diversification distinguishable, and thus $\mathrm{HDD}_{H_{HP}}(G)=1$. A detailed proof can be found in Sec.VII-A.
@@
Hard masking imposes a strict rule: any connection that has appeared at a previous hop is excluded from later hops. While this prevents redundant aggregation, it can also remove structurally important edges that act as information bridges across different regions. Such edges often participate in multiple short paths and facilitate signal propagation between cohesive substructures. To address this limitation, we introduce an adaptive masking scheme that selectively preserves connections whose multi-hop strength increases relative to the previous hop.
@@
Why the strengthened edges matter. Among existing edges, strengthened edges from previous hops are precisely the ones supported by many (preferentially low-degree) common neighbors, making them highly embedded and triangle-rich. Adaptive masking inherently favors edges that participate in dense triadic structures, which serve as stable relational backbones in local communities. This intuition can be formalized through the following theorem, which provides a lower bound on the triangle support of adaptively retained edges.
@@
A detailed proof can be found in Sec.VII-B.
Hence, edges that survive as strengthened paths under adaptive masking participate in a provably large number of triangles locally, yielding a compact, triangle-rich backbone that supplies robust seeds for community search.
@@
$\alpha_{ij}^{(k)}=\sigma\!\bigl((W\mathbf{h}_{i})^{\!\top}(W\mathbf{h}_{j})\bigr),\quad\tilde{A}^{(k)}=\hat{A}^{(k)}\odot\alpha^{(k)},$ (6)
where $W$ is a linear projection, $\mathbf{h}_{i}$ denotes the node embedding, and $\sigma$ is the sigmoid function. This local normalization allows each node to emphasize semantically consistent neighbors and mitigate noisy ones, ensuring stable propagation and balanced gradient flow under strong heterophily.
@@ IV-C Frequency Channels
Following ACM [ACM_2022], we utilize frequency channels to process both homophilic and heterophilic patterns in the graph. However, unlike the original ACM, which recursively propagates features across layers, our framework leverages hop-distinct channels, so there is no recursive propagation from previous layers. Each hop channel is computed independently, as detailed in previous sections.
To avoid the prohibitive cost of constructing hop-wise Laplacian matrices, we exploit the fact that for any normalized adjacency $A$ and corresponding Laplacian L=IAL=I-A, the high-pass response for a feature matrix $X$ can be efficiently computed as LX=XAXLX=X-AX. This insight allows us to obtain high-pass signals without explicit Laplacian construction. The high-pass feature is obtained as follows:
Raw distinctive hop. For each hop $k$, with $A^{(k)}$ denoting the exact-$k$ adjacency, the high-pass output is computed as:
@@
Local weight renormalization. When applying local weight renormalization, we first reweight the hop-$k$ adjacency by edge-wise attention as in Eq. (6), and then propagate features. The low-pass feature is therefore
$X^{(k)}_{\mathrm{LP}}\;=\;\tilde{A}^{(k)}X\;=\;\bigl(\hat{A}^{(k)}\odot\alpha^{(k)}\bigr)\,X.$ (8)
The complementary high-pass branch uses the complementary edge weights before propagation:
$\tilde{A}^{(k)}_{\mathrm{HP}}\;=\;\hat{A}^{(k)}\odot\bigl(1-\alpha^{(k)}\bigr),\qquad X^{(k)}_{\mathrm{HP}}\;=\;\tilde{A}^{(k)}_{\mathrm{HP}}X.$ (9)
In summary, this design allows our model to compute both low-pass (aggregation) and high-pass (diversification) responses for each hop without redundant memory usage, ensuring frequency-aware feature extraction across the network.
@@
Linear transformation. Before mixing, channel features are projected into a unified latent space via linear transformations:
@@
where $W_{\mathrm{LP}},W_{\mathrm{HP}}\in\mathbb{R}^{d\times h}$ are learnable parameters and $d$ is the input size, $h$ is the hidden size. This ensures both low-frequency and high-frequency channels are brought to a consistent feature space, while enhancing model expressiveness.
@@
Hop channel mixing via attention bank. Another approach to fuse hop features is via a per-class attention bank. Given a learnable class bank $W_{\mathrm{bank}}\in\mathbb{R}^{d\times c}$, we first compute a node-wise class attention, and derive the adaptive feature weight:
$w_{\mathrm{class}}=\alpha_{\mathrm{class}}W_{\mathrm{bank}}^{\top},\quad\alpha_{\mathrm{class}}=\mathrm{softmax}(H^{(0)}W_{\mathrm{bank}}).$ (15)
We first apply $w_{\mathrm{class}}$ to scale features according to the specified class pattern, then utilize the same class specified weight for the computation of hop-wise weights $\alpha_{\mathrm{hop}}$:
$H_{\mathrm{final}}=\sum_{k=1}^{K}\alpha_{\mathrm{hop},k}\cdot\left(H^{(k)}\odot w_{\mathrm{class}}\right),$ (16)
where $\odot$ denotes element-wise multiplication. This procedure adaptively weights features and hop messages via attention.
Notably, the filter bank fusion approach aligns naturally with our theoretical analysis in Theorem 1. By leveraging class-aware feature weighting, this strategy enables strict filtering of irrelevant signals and is critical for establishing high-order diversification distinguishability, as shown in our proof. Overall, this 2D-channel mixing strategy integrates both frequency and hop-specific information, enhancing the representational capacity of the model for AdaptCS.
@@ V Efficient Optimization
Although distinctive-hop channels successfully disentangle multi-hop semantics, explicitly computing exact-hop adjacency matrices $\hat{A}^{(k)}$ introduces severe memory limitations, especially on large or dense graphs. Unlike standard multi-hop methods such as SGC [sgc_2019] and MixHop [Mixhop], which efficiently precompute high-order features using iterative operations like $A(AX)$ and avoid ever materializing high-order adjacency matrices, our approach requires explicit computation of terms such as (AAA)X(AA-A)X to isolate purely two-hop neighbors. However, forming $AA$ results in significant memory overhead even when $A$ is sparse. This explicit construction of high-order adjacency matrices rapidly becomes the primary bottleneck, causing out-of-memory (OOM) errors for even small values of $k$, and fundamentally restricting the scalability of distinctive-hop channels on very large or dense datasets.
To address this critical bottleneck, we introduce a memory-efficient optimization using low-rank singular value decomposition (SVD) to approximate all distinctive-hop computations directly in a compressed latent subspace, eliminating the need to construct or store high-order adjacency matrices explicitly.
@@ V-A Low-rank distinctive-hop computation via SVD
To eliminate explicit high-order adjacency matrices, we adopt a rank-$r$ singular-value decomposition (SVD) approximation of the normalized one-hop adjacency operator. Specifically, we factorize the symmetrically normalized adjacency (low-pass channel) as $\hat{A}=U\Sigma V^{\!\top},\quad\text{where}\quad U,V\in\mathbb{R}^{n\times r},\;\Sigma=\operatorname{diag}(s_{1},\dots,s_{r}).$ Consequently, for any hop $k\geq 1$, we obtain the approximated matrix as
$\hat{A}^{k}=U\Sigma^{k}V^{\!\top},\quad\text{with}\quad\Sigma^{k}\coloneq\operatorname{diag}\!\bigl(s_{1}^{k},\dots,s_{r}^{k}\bigr).$ (17)
Therefore, the difference between consecutive adjacency powers, defining the distinctive-hop adjacency, becomes
$\hat{A}^{(k)}\;=\;\hat{A}^{k}-\hat{A}^{k-1}\;=\;U\bigl(\Sigma^{k}-\Sigma^{k-1}\bigr)V^{\!\top}.$ (18)
Using this representation, the low-pass feature at hop $k$ is computed entirely in the compressed subspace as
$X^{(k)}=\hat{A}^{(k)}X=\;U\Delta\Sigma^{(k)}\bigl(V^{\!\top}X\bigr),$ (19)
where the projection $V^{\!\top}X\in\mathbb{R}^{r\times d}$ is computed only once and shared across all hops, scaled separately for each hop using the diagonal $\Delta\Sigma^{(k)}$, and multiplied by $U$ for reconstruction.
This SVD-based approach yields two major advantages: First, all distinctive-hop computations are performed in the low-rank latent space, so there is no need to explicitly construct or store any $n\times n$ high-order adjacency matrices. This completely removes the memory bottleneck and enables our model to scale to graphs with hundreds of millions of edges using only a single GPU. Second, the computational process is highly efficient: only a single SVD decomposition and one feature projection ($V^{\top}X$) are required. All subsequent hop-specific computations are reduced to fast diagonal matrix multiplications in the small $r$-dimensional space, which greatly accelerates the overall computation and makes the method suitable for large-scale and high-order graph analysis.

### v2

@@ IV Methodology Phase i: Graph Encoding
Adaptive Community Search (AdaptCS) is a two-phase framework consisting of graph encoding and online search. During the graph encoding phase, AdaptCS learns community-oriented node representations. In the online search phase, the model leverages these learned representations to retrieve a meaningful query-centric community given an input query. This section presents the technical details of the AdaptCS encoder; the online search algorithm is described in Sec. VI.
@@ IV-A Encoder Frameworks
Figure 2 gives an end-to-end overview of our AdaptCS encoder. Starting from an input graph $G=(V,E,X)$, the model framework disentangles hop distance (Distinctive-hop channels) and spectral frequency (Frequency channels), then fuses the resulting 2D signal by a channel mixer.
Stage 1: Distinctive-hop channels (Sec.IV-B). In Figure 2(a), given a node $v_{i}$, we recursively conduct $k$-hop neighborhoods aggregation to construct distance-aware features. The $k$-th channel only aggregates information from neighbors with an exact shortest distance of $k$, which eliminates redundancy and noisy messages during convolution. We then apply weight renormalization by assigning edge-wise similarity scores via dot products between the embedding matrix $H\in\mathbb{R}^{n\times h}$, which rescales the aggregated messages and stabilizes magnitudes.
Stage 2: Frequency channels (Sec.IV-C). Frequency filters (low pass and high pass) process the hop-specific feature matrices to produce two complementary views: a smooth (homophilic) representation and a non-smooth (heterophilic) representation (Figure 2(b)). This separation allows each channel to specialize in either cohesive similarity or distinctive contrast without mutual interference.
Stage 3: 2D-channel mixing (Sec.IV-D). In Figure 2(c), we first use attention to fuse the low- and high-frequency channels within each hop, and then concatenate the fused representations across all hops. Next, a multi-layer perceptron (MLP) or a per-class attention bank is applied, producing a unified node embedding that integrates both local details and long-range context in a compact representation.
@@ IV-B Distinctive Hop Channels (Distance-aware)
In this section, we propose distinctive hop channels to explicitly process signals from different hops of neighbors, designed for different homophily levels. Distinctive hop channels address three limitations in traditional multi-hop aggregation approaches. First, this approach solves the flip effect, where signals from each hop are processed independently, handling the complex semantics between higher-order relations. Second, the distinctive hop adjacency matrices are shortest-distance aware, where each node in the $k$-th hop channel accurately represents edges with the shortest path distance of exactly $k$. Third, this approach reduces redundancy and improves space efficiency since the sum of all distinctive hop adjacency matrices precisely reconstructs the cumulative adjacency $\hat{A}^{k}$, avoiding repetitive edge representation. To explicitly measure whether a representation preserves label-consistent similarity under such settings, we introduce the following notion.
Definition IV.1 (Heterophily-aware Node Distinguishability (HND)).
Given a labeled graph $G=(V,E)$, let $z_{u}$ denote the class label of node $u\in V$. Let $\{h_{u}\}_{u\in V}$ be the hidden state and define similarity $S(u,v)=\mathrm{sim}(h_{u},h_{v})$. In multi-class heterophily ($c>2$), node $u$ is heterophily-distinguishable if
$\mathbb{E}\!\left[S(u,v)\mid z_{v}=z_{u}\right]\ \geq\ \mathbb{E}\!\left[S(u,v)\mid z_{v}\neq z_{u}\right].$
We define $\mathrm{HND}(G)$ as the fraction of nodes $u$ for which the above condition is well-defined and holds.
HND serves as a direct representation-level criterion: a higher value indicates that the learned embedding geometry remains consistent with labels despite heterophilic connectivity. Our key idea is to make multi-hop aggregation distance-aware by constructing exact-$k$-hop channels, and then use a prototype bank to select which hop signals should dominate the final representation. The following proposition summarizes this mechanism and its implications for HND.
Proposition 1 (Distance-aware hop selection improves HND).
AdaptCS constructs exact-$k$-hop representations $\{h_{u}^{(k)}\}$ and fuses them via prototype-gated hop attention. Under hop separability and mild gating margins, the fusion upweights the label-consistent hop, yielding embeddings that satisfy the HND ordering in Definition IV.1 and thus improve $\mathrm{HND}(G)$. The formal guarantee is given in Theorem 2.
@@
Hard masking imposes a strict rule: any connection that has appeared at a previous hop is excluded from later hops. While hard masking avoids redundant aggregation, it can also discard structural signals in densely-connected regions, where the same relations are repeatedly reinforced across multiple hops. Hence, higher-hop propagation is forced to rely on newly emerging connections, which are typically less stable and more noise-sensitive. To address this limitation, we introduce an adaptive masking scheme that selectively preserves edges whose multi-hop strength increases from the previous hop.
@@
Why the strengthened edges matter. At k=2k{=}2, strengthened existing edges $(u,v)\in E$ (i.e., $(\hat{A}^{2}-\hat{A})_{uv}>0$) are exactly those receiving substantial two-hop support from common neighbors, hence they are highly embedded and triangle-rich.
Why does this benefit the community search. (i) Triangle-rich edges anchor locally dense subgraphs, concentrating similarity and making the endpoints more likely to fall into the same community; (ii) multiple common neighbors create more two-hop paths that average out label noise, bringing endpoints representation closer together even under heterophily (the neighbors’ labels may vary, but the shared-neighborhood signal remains coherent). This yields a concrete, verifiable advantage at two hops, formalized below.
@@
A detailed proof appears in Appendix A online [39].
Beyond k=2k{=}2. While Theorem 1 specifically certifies local triadic density at two hops, the same edges typically retain larger multi-hop mass at higher orders (via degree-normalized walk contributions), making such node pairs more likely to reside in the same locally dense subgraph.
@@
$\alpha_{ij}^{(k)}=\sigma\!\bigl((W\mathbf{h}_{i})^{\!\top}(W\mathbf{h}_{j})\bigr),\ \tilde{A}^{(k)}=RN(\hat{A}^{(k)}\odot\alpha^{(k)}),$ (6)
where $W$ is learnable parameters, $\mathbf{h}_{i}$ denotes the node embedding for node $v_{i}$, $\sigma$ is the sigmoid function, $\alpha^{(k)}\in\mathbb{R}^{n\times n}$ is the weights, $RN$ is RowNorm which rescales each row to sum to one. This local normalization enables each node to prioritize semantically consistent neighbors and minimize the influence of noisy ones, thereby ensuring stable propagation and balanced gradient flow under heterophily.
@@ IV-C Frequency Channels
In the previous sections, we have constructed hop-distinct propagation operators $\{\hat{A}^{(k)}\}$, which naturally serve as low-pass channels that aggregate information from the exact-$k$ neighborhood. Following ACM [28], we further incorporate a high-pass channel to capture heterophily-informative components that low-pass aggregation may over-mix. Unlike the original ACM, which recursively propagates features across layers, our framework leverages hop-distinct channels.
To avoid the prohibitive cost of explicitly constructing hop-wise Laplacian matrices, we exploit the identity that for any normalized adjacency $A$ with Laplacian L=IAL=I-A, the high-pass response can be computed efficiently as LX=XAXLX=X-AX.
Raw distinctive hop. For each hop $k$, with $A^{(k)}$ denoting the exact-$k$ adjacency, the hop-wise Laplacian operator is:
@@
Local weight renormalization. When applying local weight renormalization, we first reweight the hop-$k$ operator by edge-wise attention as in Eq. (6) and then propagate features. The $k$-hop features aggregated by a distinct low-pass filter $\tilde{A}^{(k)}$ is
$X^{(k)}_{\mathrm{LP}}\;=\;\tilde{A}^{(k)}X\;=\;RN\bigl(\hat{A}^{(k)}\odot\alpha^{(k)}\bigr)\,X.$ (8)
The reweighted high-pass branch is as follows:
$\tilde{A}^{(k)}_{\mathrm{HP}}\;=\;RN\bigl(\hat{A}^{(k)}\odot(1-\alpha^{(k)})\bigr),$ (9)
$X^{(k)}_{\mathrm{HP}}\;=\;\tilde{A}^{(k)}_{\mathrm{HP}}X.$ (10)
In summary, our hop-distinct design computes both low-pass and high-pass responses per hop using shared operators, avoiding explicit Laplacian construction and additional memory overhead while enabling frequency-aware feature extraction.
@@
Linear transformation. Before mixing, channel features are projected into a unified latent space via linear transformations, $H_{\mathrm{LP}}^{(k)}$ and $H_{\mathrm{HP}}^{(k)}$ are hidden state aggregated by low/high pass filter at the $k$-hop, calculate as follows:
@@
where $W_{\mathrm{LP}},W_{\mathrm{HP}}\in\mathbb{R}^{d\times h}$ are learnable parameters and $d$ is the input size, $h$ is the hidden size.
@@
Hop channel mixing via attention bank. Another approach to fuse hop features is via a per-class attention bank. Given a learnable class bank $P\in\mathbb{R}^{d\times c}$, we first compute a node-wise class attention, and derive the adaptive feature weight:
$w^{\mathrm{class}}=\alpha^{\mathrm{class}}P^{\top},\quad\alpha^{\mathrm{class}}=\mathrm{softmax}(H^{(0)}P).$ (16)
We first apply $w^{\mathrm{class}}$ to scale features according to the specified class pattern, then utilize the same class specified weight for the computation of hop-wise weights $\alpha^{\mathrm{hop}}$:
$H_{\mathrm{final}}=\sum_{k=1}^{K}\alpha^{\mathrm{hop}}_{k}\cdot\left(H^{(k)}\odot w^{\mathrm{class}}\right),$ (17)
where $\odot$ denotes element-wise multiplication. This procedure adaptively weights features and hop messages via attention. Notably, the filter bank fusion approach aligns naturally with our theoretical analysis in Theorem 2.
Negative Log-Likelihood Loss. Given labels $y_{i}\in\{1,\dots,C\}$, the negative log-likelihood loss is
$\mathcal{L}_{\mathrm{NLL}}=-\frac{1}{N}\sum_{i=1}^{N}\log\Big(\textit{softmax}(H_{\mathrm{final}}[i,:])_{y_{i}}\Big).$ (18)
It encourages high probability on the ground-truth class by penalizing low confidence on $y_{i}$.
IV-E Extension Discussion
While AdaptCS is evaluated on undirected graphs with non-overlapping communities, the proposed model can be extended using standard techniques from prior work [40, 33].
Directed graphs. Directed graphs can be handled mainly in the encoder by adopting standard directional message passing, e.g., separating in/out neighborhoods, followed by a fusion of these channels. This is consistent with SDGNN [40], which is guided by status/balance theory and learns representations by jointly modeling link direction, link sign, and signed directed triads (triangles) as auxiliary supervision.
Graphs with labeled edges. Graphs with labeled (or typed) edges can be supported by incorporating edge semantics into the encoder, e.g., using relation-specific transformations or edge-aware attention to modulate message passing for different edge types. In this way, the learned representations reflect both structural connectivity and edge-level semantics, while the online search procedure remains unchanged.
Overlapping communities. Our framework can be extended to overlapping community search by adopting the Sparse Subspace Filter (SSF) paradigm in SMN [33]. This subspace-based retrieval can be integrated into our offline–online pipeline with minimal changes, while preserving efficient online search.
@@ V Efficient Optimization
Although distinctive-hop channels successfully disentangle multi-hop semantics, explicitly computing exact-hop adjacency matrices $\hat{A}^{(k)}$ introduces severe memory limitations, especially on large or dense graphs. Unlike standard multi-hop methods such as SGC [41] and MixHop [36], which efficiently precompute high-order features using iterative operations like $A(AX)$ and avoid ever materializing high-order adjacency matrices, our approach requires explicit computation of terms such as (AAA)X(AA-A)X to isolate purely two-hop neighbors. However, forming $AA$ results in significant memory overhead even when $A$ is sparse. To address this critical bottleneck, we introduce a memory-efficient optimization using low-rank singular value decomposition (SVD) to approximate distinctive-hop computations in a compressed latent subspace, eliminating the need to construct high-order adjacency matrices explicitly.
@@ V-A Low-rank distinctive-hop computation via SVD
To avoid explicit high-order adjacency matrices, we adopt a rank-$r$ truncated SVD approximation of the normalized one-hop adjacency operator. Specifically, we factorize the symmetrically normalized adjacency (low-pass channel) as $\hat{A}\approx U_{r}\Sigma_{r}V_{r}^{\!\top},\quad\text{where}\quad U_{r},V_{r}\in\mathbb{R}^{n\times r},\;\Sigma_{r}=\operatorname{diag}(s_{1},\dots,s_{r}).$ For any hop $k\geq 1$, the approximated matrix is as
$\hat{A}^{k}\approx U_{r}\Sigma_{r}^{k}V_{r}^{\!\top},\quad\text{with}\quad\Sigma_{r}^{k}\coloneq\operatorname{diag}\!\bigl(s_{1}^{k},\dots,s_{r}^{k}\bigr).$ (19)
Therefore, the distinctive-hop adjacency becomes
$\hat{A}^{(k)}=\hat{A}^{k}-\hat{A}^{k-1}\approx U_{r}\bigl(\Sigma_{r}^{k}-\Sigma_{r}^{k-1}\bigr)V_{r}^{\!\top}.$ (20)
Using this representation, the low-pass feature at hop $k$ is computed in the compressed subspace as
$X^{(k)}=\hat{A}^{(k)}X\approx U_{r}\Delta\Sigma_{r}^{(k)}\bigl(V_{r}^{\!\top}X\bigr),$ (21)
where $V_{r}^{\!\top}X\in\mathbb{R}^{r\times d}$ is computed once and shared across hops.
This SVD-based approach yields two major advantages: First, all distinctive-hop computations are performed in the low-rank latent space, so there is no need to explicitly construct or store any $n\times n$ high-order adjacency matrices. This completely removes the memory bottleneck and enables our model to scale to graphs with hundreds of millions of edges using only a single GPU. Second, the computational process is highly efficient: only a single SVD decomposition and one feature projection ($V_{r}^{\top}X$) are required. All subsequent hop-specific computations are reduced to fast diagonal matrix multiplications in the small $r$-dimensional space, which greatly accelerates the overall computation and makes the method suitable for large-scale and high-order graph analysis.